OpenStudy (anonymous):

Simplify the expression. -4+6x(-1)

OpenStudy (anonymous):

=-4-6x

rvc (rvc):

6x(-1)=?

OpenStudy (anonymous):

A. –10 B. –2 C. 2 D. 10

rvc (rvc):

-10

OpenStudy (anonymous):

just plug in x=-1 you will get=-4-4=-10

OpenStudy (anonymous):

order of operations. multiplty 6x by -1 then do -4 + ..... when u plug in x=-1 = -10